OpenStudy (anonymous):

find the slope of the curve y= sin x/x at the point x=π/4.

OpenStudy (anonymous):

Find dy/dx at x = pi/4 Use quotient rule to differentiate

OpenStudy (anonymous):

dy/dx=(cosx/x)-(sinx/x^2) dy/dx|_{pi/4}=((4/pi)-16/pi^2)/sqrt(2)

OpenStudy (anonymous):

what did u do in the last step?

OpenStudy (anonymous):

i just put x=pi/4

OpenStudy (anonymous):

dy/dx|_{pi/4} means value of dy/dx at pi/4

OpenStudy (anonymous):

oh ok :)